OpenStudy (anonymous):

multiplicative inverse of \[2+3i \div2-3i\]

myininaya (myininaya):

multiplicative inverse of \[\frac{2+3i}{2-3i}\] is \[\frac{2-3i}{2+3i}\]

myininaya (myininaya):

since \[\frac{2+3i}{2-3i} \cdot \frac{2-3i}{2+3i}=1\]

OpenStudy (anonymous):

is there a further step because thats none of the answers match that

myininaya (myininaya):

you can always right it in a+bi form

OpenStudy (anonymous):

the all have like 5-12i and tings of that sort

myininaya (myininaya):

just do \[\frac{2-3i}{2+3i} \cdot \frac{2-3i}{2-3i}\]

OpenStudy (anonymous):

they*